Hillary's Lexus travels 30 mph faster than Bill's Harley. In the same time that Bill travels 75 mi, Hillary travels 120 mi. Find their speeds.
Answer by scott8148(6628)

"Hillary's Lexus travels 30 mph faster than Bill's Harley" ___ HL = BH +30

time = distance / rate

"In the same time that Bill travels 75 mi, Hillary travels 120 mi" ___ 75 / BH = 120 / HL

"cross" multiplying ___ 75 HL = 120 BH

substituting ___ 75 (BH + 30) = 120 BH ___ 75 BH + 2250 = 120 BH ___ 2250 = 45 BH ___ 50 = BH

substituting ___ HL = 50 + 30 ___ HL = 80